# A General Error-Theoretical Analysis Framework for Constructing Compression Strategies

### v1

@@ 1 Introduction
Against this backdrop, mixed compression techniques have become a mainstream approach for model compression. Due to the varying contributions of different layers to overall model performance, different layers exhibit significantly different tolerance levels to compression. Adopting layer-wise differentiated compression strategies can thus maximize compression rates while minimizing performance degradation. For example, in low-rank decomposition, different layers have varying ranks. In quantization, some critical layers require higher precision, while others can use lower precision. However, the main challenge with this approach lies in the exponentially growing search space for determining the optimal mixed-precision quantization configuration. Specifically, for a neural network with layers, where each layer can select from four possible bit-widths (e.g., 2/3/4/8 bits), the search space grows to . The huge search space makes it almost impossible to find the optimal configuration that can maintain good generalization performance and meet hardware efficiency.
Several representative compression methods have been proposed to address these challenges. For instance, DMBQ (Zhao et al., 2021) pre-emptively searches for the optimal bit-width configuration in the distribution space and dynamically selects compression settings during training. The HAWQ (Dong et al., 2019b) series computes layer-wise Hessian information to assess the relative sensitivity of each layer, thereby determining the optimal compression configuration. However, these approaches suffer from the following limitations: (1) Lack of optimality explanation: Existing compression configurations often rely on heuristic approaches, lacking a clear theoretical foundation to justify their optimality. (2) Neglect of error correlation: These methods typically decouple the optimization of model performance error and compression error, failing to systematically analyze their theoretical interdependence. (3) Limited adaptability to aggressive quantization: At lower bit widths, these methods struggle to handle degradation caused by compression errors, which can disrupt retraining and result in significant performance drops.
@@ 4.1 Datasets and Details
ResNet-50 | ABC-Net | 76.10 | 5 | 5 | 6.40 | 6.40 | 70.10 | -6.00 |
LQ-Nets | 76.40 | 3 | 3 | 5.99 | 10.64 | 74.20 | -2.20 | |
LQ-Nets | 76.40 | 4 | 4 | 5.11 | 7.99 | 75.10 | -1.30 | |
DoReFa | 76.90 | 4 | 4 | 5.11 | 7.99 | 71.40 | -5.50 | |
@@
4.2 Ablation
Differential Expansion. Differential expansion requires ensuring that the neighborhood around the expansion point is sufficiently small to guarantee that the theoretical approximation effectively represents the actual values within this neighborhood. Figure 2 shows the discrepancies between theoretical and actual values under different perturbation errors introduced into each layer. Taking ResNet-18 as an example, the gap between theoretical and actual values is minimal across layers, which validates the effectiveness of CET. Additionally, error perturbation experiments reveal that different layers exhibit distinct levels of error tolerance.
@@ 4.3 Comparison
Algorithm Accuracy. As shown in Table 1, we conducted a comprehensive comparison between CET and existing quantization methods (Zhang et al., 2018; Lin et al., 2017; Zhou et al., 2016; Choi et al., 2018; Han et al., 2015; Wang et al., 2019; Dong et al., 2019a). For ResNet-18 (He et al., 2016), CET achieved nearly lossless weight quantization, with model performance degrading by only 0.01%, while significantly reducing the model size, achieving a compression ratio of over 11. This demonstrates the effectiveness of CET’s geometric analysis based on second-order information, enabling precise weight compression along the long-axis direction and selecting appropriate error tolerance for each layer. Furthermore, to investigate the impact of activation compression on weight quantization, we performed quantization on activations with different bit widths (8-bit and 4-bit). The results show that CET can maximize model compression while maintaining accuracy. Notably, CET does not rely on fine-tuning but directly determines the optimal bit-width allocation, which distinguishes it from methods like MCKP that require fine-tuning.
@@ 4.4 Discussion
Limitations. Ideally, the CET framework requires precise Hessian information to ensure accurate analysis of quantization errors. However, since directly computing the full Hessian matrix is infeasible in practice, we approximate it using the Lanzcos algorithm. While this approximation introduces some degree of error, our perturbation analysis of the Lanzcos algorithm reveals that it has a strong error tolerance. The approximate Hessian information obtained is robust in most cases, ensuring the practical applicability of the CET framework.

### v2

@@ Abstract
zhangby01@pcl.ac.cn, {chengdaning, zyq}@ict.ac.cn,
tumeiqi24@connect.hku.hk, fangminghk@gmail.com, mijiake@mail.scut.edu.cn
@@ 1 Introduction
Mixed compression techniques have become a mainstream approach for model compression. Due to the varying contributions of different layers to overall model performance, compared with the unified compression approach, different layers exhibit significantly different tolerance levels to compression. Adopting layer-wise differentiated compression strategies can thus maximize compression rates while minimizing performance degradation. For example, in low-rank decomposition, different layers have varying ranks. In quantization, some critical layers require higher precision, while others can use lower precision. However, the main challenge with this approach lies in the exponentially growing search space for determining the optimal mixed-precision quantization configuration. Specifically, for a neural network with layers, where each layer can select from four possible bit-widths (e.g., 2/3/4/8 bits), the search space grows to . The huge search space makes it almost impossible to find the optimal configuration that can maintain good generalization performance and meet hardware efficiency.
Several representative compression methods (Zhao et al., 2021; frantar2022gptq; Dong et al., 2019a) have been proposed to address these challenges. For instance, DMBQ pre-emptively searches for the optimal bit-width configuration in the distribution space and dynamically selects compression settings during training. The HAWQ series computes layer-wise Hessian information to assess the relative sensitivity of each layer, thereby determining the optimal compression configuration. However, these approaches suffer from the following limitations: (1) Lack of optimality explanation: Existing compression configurations often rely on heuristic approaches, lacking a clear theoretical foundation to justify their optimality. (2) Neglect of error correlation: These methods typically decouple the optimization of model performance error and compression error, failing to systematically analyze their theoretical interdependence. (3) Limited adaptability to aggressive quantization: At lower bit widths, these methods struggle to handle degradation caused by compression errors, which can disrupt retraining and result in significant performance drops.
@@ 4.1 Datasets and Details
| ResNet-50 | DoReFa | 76.90 | 4 | 4 | 5.11 | 7.99 | 71.40 | -5.50 |
@@ 4.2 Comparison
Algorithm Accuracy. As shown in Table 1, we conducted a comprehensive comparison between CET and existing quantization methods (Lin et al., 2017; Zhang et al., 2018; Zhou et al., 2016; Choi et al., 2018; Han et al., 2015; huang2025hessian; xu2024ptmq; frantar2022optimal). For ResNet-18 (He et al., 2016), CET achieved nearly lossless weight quantization, with model performance degrading by only 0.01%, while significantly reducing the model size, achieving a compression ratio of over 11. This demonstrates the effectiveness of CET’s geometric analysis based on second-order information, enabling precise weight compression along the long-axis direction and selecting appropriate error tolerance for each layer. Furthermore, to investigate the impact of activation compression on weight quantization, we performed quantization on activations with different bit widths (8-bit and 4-bit). The results show that CET can maximize model compression while maintaining accuracy. Notably, CET does not rely on fine-tuning but directly determines the optimal bit-width allocation, which distinguishes it from methods like MCKP that require fine-tuning.
@@
4.3 Ablation
Differential Expansion. Differential expansion ablation is to explore the effective neighborhood of CET at each layer. Differential expansion requires ensuring that the neighborhood around the expansion point is sufficiently small to guarantee that the theoretical approximation effectively represents the actual values within this neighborhood. Figure 2 shows the discrepancies between theoretical and actual values under different perturbation errors introduced into each layer. Taking ResNet-18 as an example, the gap between theoretical and actual values is minimal across layers, which validates the effectiveness of CET. Additionally, error perturbation experiments reveal that different layers exhibit distinct levels of error tolerance.
@@ 4.4 Discussion
Limitations. Ideally, the CET requires precise Hessian information to ensure accurate analysis of quantization errors. However, since directly computing the full Hessian matrix is infeasible in practice, we approximate it using the Lanzcos algorithm. While this approximation introduces some degree of error, our perturbation analysis of the Lanzcos algorithm reveals that it has a strong error tolerance. The approximate Hessian information obtained is robust in most cases, ensuring the practical applicability of the CET.